OpenStudy (anonymous):

Vivian combined different amounts of white paint, blue paint, and green paint to make 145 milliliters (ml) of paint of a new shade. The amount of white paint was x ml. The amount of blue paint was 10 ml less than twice the amount of white paint, and the amount of green paint was half the amount of blue paint. This is represented in the equation below. x + (2x – 10) + one half (2x – 10) = 145 What was the difference in the amounts of white paint and blue paint Vivian combined? PLEASE HELP !!!!!!!!!!!!!

OpenStudy (kropot72):

To solve the equation for x you need to follow these steps: Remove the brackets by multiplying out. Group like terms. Can you do this?

OpenStudy (anonymous):

i'm studying for a test on monday and this is the only one that confused me @kropot72

OpenStudy (kropot72):

To remove the brackets you need to do this: \[2x \times \frac{1}{2}-10\times \frac{1}{2}=?\]

OpenStudy (anonymous):

then solve the equation? @kropot72

OpenStudy (kropot72):

Do you have a problem with removing the brackets?

OpenStudy (anonymous):

x + (2x − 10) +1/2 (2x − 10) = 145 x + 2x - 10 + x - 5 = 145 4x - 15 = 145 4x = 145 + 15 4x = 160 x = 40 ml @kropot72 am i correct?

OpenStudy (kropot72):

Good work! Now you need to find the amount of blue paint which is \[(2\times 40)-10=?\ ml\] Then subtract the smaller amount from the larger amount of the two paints to find the answer.

OpenStudy (anonymous):

equals 70 ml @kropot72

OpenStudy (anonymous):

then i would subtract 70-40 and my final answer would be 30 ml @kropot72

OpenStudy (kropot72):

Yes, your final answer is correct. Good work :)

OpenStudy (anonymous):

Solve for y: 2 over 3y + 4 = 6. @kropot72 should simplify 2 over 3y first or ??